# Measuring Star Formation Histories from Asymptotic Giant Branch Stars: A Demonstration in M31

### v2

@@ 8 Conclusions
There are several advantages to measuring SFHs from resolved AGB stars in the NIR. First, AGB stars are much brighter () than optical age sensitive features (e.g., red clump, MSTO). AGB star in the NIR are also less impacted by dust. In the NIR AGB stars occupy a distinct region of the CMD with little-to-no overlap with other populations (e.g., RGB stars, MW foreground). In the optical, AGB stars are challenging to separate from other phases of evolution that are not age sensitive (e.g., the RGB) compromising their utility for SFH work.
@@
We use only AGB stars from ground-based UKIRT NIR photometry to measure the SFH over the past 8 Gyr of M31’s outer disk covered by PHAT. We find our SFH to be in excellent agreement with the area-matched PHAT SFH from W17, which was derived from an optical HST-based CMD. The lookback time of this comparison is set by the fidelity of the PHAT SFH; the AGB star CMDs shows populations that are older than 8 Gyr. Our tests show that AGB stars-based SFHs can be recovered to ages older than 8 Gyr ago.
@@
By sub-sampling the number of AGB stars used for comparison against the PHAT SFH, we show that even a few hundred AGB stars can provide a robust SFH recovery. We attribute this to the strong age-sensitivity of AGB stars in the NIR.
As a practical illustration of our AGB star SFH methodology, we measure the spatially resolved SFH of M31’s inner stellar halo ( kpc. We find a nearly global burst of star formation Gyr ago and much lower level, spatially distributed star formation from Gyr ago. Regions associated with NGC 205 and M32 appear to the most active. We find a total of formed in M31’s inner stellar halo in the last 8 Gyr.
We use synthetic data to show that (a) SFHs only using NIR observations of AGB stars can robustly be recovered to lookback times of Gyr ago and (b) JWST CMDs of AGB stars (e.g., ) can be used to recovery SFH across all cosmic time with similar fidelity.
@@
Appendix A How many stars are needed for an accurate AGB SFH measurement?
In this section, we investigate the effect of decreased sample size on the precision and accuracy of our AGB SFH measurements. As described in §5.2.2, a sample of 23,000 AGB stars delivered small statistical uncertainties that agree well systematically with the measurement from W17. To test how the uncertainties and bias change as a function of sample size, first we randomly sample 11500 AGB stars, or about 1/2 of our main sample, and re-measure the AGB SFH. We repeat this test eight more times until our sample size decreases to AGB stars. In Figure 16, we show three representative fits measured from random samples of 5900 AGB stars, 760 AGB stars, and 280 AGB stars.
As expected, the statistical uncertainties increase with decreased sample size. We note a sample of 700 AGB stars deliver an AGB SFH closely matching (albeit with larger statistical errors) the AGB SFH measured from the full sample. For a sample of 300 AGB stars, the measured AGB SFH begins to systematically deviate slightly from the AGB measured from 23,000 AGB stars, particularly at older ages. Therefore, we conclude that accurate AGB SFH measurements should be based on samples of at least AGB stars.
Appendix B Photometric Quality-metric cuts
@@
We clean the UKIRT photometry used in the main text via the photometric uncertainty as a function of their J-band and K-band magnitudes. These cuts were also used in Lee (2023) who used this same photometric catalog to measure the distance to M31 with AGB stars. The functional forms of these functions are shown below and also plotted in Figure B1:
Appendix C Recovery of Oldest AGB Stars
Throughout the paper, we limit our simulated and real SFHs to a lookback time of Gyr ago. This age was motivated by limitations of the SFHs presented by PHAT. As discussed in W17, they only model one age bin between 8 and 14 Gyr ago, owing to the modest depth of their data. The utility of this age bin is limited for SFH validation as it essentially provides a total mass and nothing else. We found it far more valuable to focus on both comparing the normalization and timing of star formation between the our NIR AGB star approach and the SFHs from PHAT optical CMDs. When we find a suitable dataset, i.e., that has enough AGB stars and an optical CMD that reaches fainter than the oldest MSTO, we will undertake a similar SFH comparison to older ages.
@@ Appendix C Recovery of Oldest AGB Stars
In the meantime, we use simulated data to show that, in principle, AGB star-based SFHs can be recovered back to the oldest epochs of star formation. The set up of the mock tests are the same as described in §5.1. The results for a constant SFH in the filter combination are shown in Figure C1. The mock tests show excellent recovery of the input SFH.
@@
In Figure C2, we undertake the same test only using an example JWST filter set . The motivation for this filter combination is that it is particularly efficient for measuring precise TRGB distances with JWST (e.g., Newman et al., 2024). Such data would also well-suited for measuring AGB star-based SFHs. As with the other simulated data experiments, the recovered SFH is in good agreement with the input SFH. Deviations in the SFR at early time are not statistically significant and are similar to fluctuations in CMD-based SFH recovery discussed in Dolphin (2002), and further emphasize the importance of robust statistical uncertainties (e.g., Dolphin, 2013). Our purpose was largely to demonstrate the utility of excellent JWST TRGB data for this additional use case. We plan a more exhaustive exploration of JWST (as well as Roman and Euclid) filter combinations in a future paper.

### v3

@@ 8 Conclusions
There are several advantages to measuring SFHs from resolved AGB stars in the NIR. First, AGB stars are much brighter () than optical age sensitive features (e.g., red clump, MSTO). AGB stars in the NIR are also less impacted by dust. In the NIR AGB stars occupy a distinct region of the CMD with little-to-no overlap with other populations (e.g., RGB stars, MW foreground). In the optical, AGB stars are challenging to separate from other phases of evolution that are not as age sensitive (e.g., the RGB) compromising their utility for SFH work.
@@
We use only AGB stars from ground-based UKIRT NIR photometry to measure the SFH over the past 8 Gyr of M31’s NE outer disk ( kpc) covered by PHAT. We find our SFH to be in excellent agreement with the area-matched PHAT SFH from W17, which was derived from an optical HST-based CMD. The lookback time of this comparison is set by the fidelity of the PHAT SFH and the depth of our ground-based UKIRT data. Our simulations show that theoretically, AGB stars-based SFHs can be recovered to ages older than 8 Gyr ago.
@@
By sub-sampling the number of AGB stars used for comparison against the PHAT SFH, we show that AGB stars can provide a robust SFH recovery. We attribute this to the strong age-sensitivity of AGB stars in the NIR.
As a practical illustration of our AGB star SFH methodology, we measure the spatially resolved SFH of M31’s inner stellar halo ( kpc. We find a nearly global burst of star formation Gyr ago and much lower level, spatially distributed recent star formation 1 – 3 Gyr ago. We find a total of formed in M31’s inner stellar halo in the last 8 Gyr.
We use synthetic data to show that (a) SFHs only using NIR observations of AGB stars can robustly be recovered to lookback times of Gyr ago and (b) JWST CMDs of AGB stars (e.g., ) can be used to recover SFH across all cosmic time with similar fidelity.
@@
Appendix A Photometric Quality-metric cuts
@@
We clean the UKIRT photometry used in the main text via the photometric uncertainty as a function of their J-band and K-band magnitudes. The functional forms of these functions are shown below and also plotted in Figure A1:
Appendix B Recovery of Oldest AGB Stars
Throughout the paper, we limit our simulated and real SFHs to a lookback time of Gyr ago. This age was motivated by the depth of our ground-based data. To verify that our AGB photometric selection criteria excluded stars with ages Gyr for our analysis in Section 5.2, we show a plot of the number of stars recovered per age bin in Figure B1. This plot demonstrates that we are recovering a negligible number of stars with ages Gyr, as expected. When we find a suitable NIR dataset, i.e., that has AGB stars and an optical CMD that reaches fainter than the oldest MSTO, we will extend the photometric selection criteria to fainter magnitudes to undertake an SFH comparison to older ages.
@@ Appendix B Recovery of Oldest AGB Stars
In the meantime, we use simulated data to show that, in principle, AGB star SFHs can be recovered back to the oldest epochs of star formation. We adopt an age range of to with a spacing of 0.1 dex, and a metallicity range of [M/H] = -1.4 to 0.0 dex with a spacing of 0.2 dex. The results for a a roughly constant star formation rate of from 100 Myr to 14.1 Gyr in the filter combination are shown in Figure B2. The mock tests show excellent recovery of the input SFH.
@@
In Figure B3, we undertake the same test only using an example JWST filter set . The motivation for this filter combination is that it is particularly efficient for measuring precise TRGB distances with JWST (e.g., Newman et al., 2024). Such data would also well-suited for measuring AGB star SFHs. As with the other simulated data experiments, the recovered SFH is in good agreement with the input SFH. Deviations in the SFR at early time are not statistically significant and are similar to fluctuations in CMD-based SFH recovery discussed in Dolphin (2002), and further emphasize the importance of robust statistical uncertainties (e.g., Dolphin, 2013). Our purpose was largely to demonstrate the utility of excellent JWST TRGB data for this additional use case. We plan a more exhaustive exploration of JWST (as well as Roman and Euclid) filter combinations in a future paper.
Appendix C AGB Age-metallicity relation
Figure C1 presents the best-fit AMR from our AGB star SFH measurements, compared with all the model fits from W17 for data kpc. We find good overall agreement between their overall shapes, mainly the noticeable drop to [M/H] dex at Gyr, albeit we recovered it at slightly earlier times, and only to [M/H] dex. We plan to further investigate how well AGB stars can recover the AMRs compared with fits derived from the oMSTO in future papers.
Appendix D How many stars are needed for an accurate AGB star SFH measurement?
In this section, we investigate the effect of decreased sample size on the precision and accuracy of our AGB star SFH measurements. As described in §5.2.2, a sample of 7700 AGB stars delivered small statistical uncertainties that agree well systematically with the measurement from W17. To test how the uncertainties and bias change as a function of sample size, first we randomly sample 3600 AGB stars, or about 1/2 of our main sample, and re-measure the AGB star SFH. We repeat this test until our sample size decreases to AGB stars. In Figure D1, we show two representative fits measured from random samples of 1000 AGB stars and 460 AGB stars.
As expected, the statistical uncertainties increase with decreased sample size. We note a sample of 1000 AGB stars deliver an AGB star SFH that agrees with (albeit with larger statistical errors) the AGB star SFH measured from the full sample, to within . For a sample of 460 AGB stars, the measured AGB star SFH begins to systematically deviate slightly from the AGB measured from 7700 AGB stars. Therefore, we conclude that accurate AGB star SFH measurements should be based on samples of at least AGB stars.